OpenStudy (anonymous):

A soccer ball is kicked with an initial speed of 8.8 m/s. After 0.179 s it is at its highest point. What was its initial direction of motion?

OpenStudy (anonymous):

Answer is in degrees (above the horizon)

OpenStudy (anonymous):

 we just need the verticle component to find the angle using inverse sine. \[V=V1 + A*T\] V1 = unknown V(at highest point) = 0 A= gravity = -9.8 m/s^2 T=.179s \[0=V1 - 9.8m/s^2 * 0.179s\] V1=1.7542m/s \[\sin \Theta = (1.7542m/s) / (8.8m/s)\] use inverse of sine to leave theta on the left alone \[\Theta =\sin^{-1} (1.75/8.8)\] \[\Theta = 11.50 degrees\]